# Is it possible to describe an electron by the evolution of a single point?

### v1

@@ Abstract
The answer to the title-question is positive. The analysis of the geometry of continuous and differentiable curves in three-dimensional Euclidean space suggests that the point represents the location of the center of charge of the electron, satisfies a system of ordinary differential equations of fourth order, and moves at the speed of light. The center of mass of the electron is a different point and will be determined by the evolution of the center of charge. It is the relative motion of the center of charge around the center of mass that gives rise to the spin and magnetic properties.
@@ 1 Introduction
The theory of continuous and differentiable curves in three-dimensional space was developped in 1847 and is known as the Frenet-Serret formalism. The points of the curve are parameterized by the arc length $s$ that represents the Euclidean distance along the curve of this point ${\bi r}(s)$ from some initial position ${\bi r}(0)$. The function ${\bi r}(s)$ and its subsequent derivatives with respect to the parameter $s$ are continuous and differentiable functions of this parameter.
@@
Along the curve, and in terms of the derivatives of the point ${\bi r}(s)$ up to the third order, it is defined a system of three orthogonal unit vectors attached to the point. They are called the tangent $\bi t$}, the normal $\bi n$} and the binormal $\bi b$}. The characteristic of any curve are the two continuous and differentiable scalar functions $\kappa(s)$ and $\tau(s)$, known as the curvature and torsion, respectively, defined at every point ${\bi r}(s)$. The curvature is expressed in terms of the first two derivatives of the function ${\bi r}(s)$ and the torsion involves the derivatives up to the third order. The knowledge of the two continuous functions $\kappa(s)$ and $\tau(s)$, for $s\in[0,L]$, together the boundary conditions ${\bi r}(0)$ and ${\bi t}(0)$, ${\bi n}(0)$ and ${\bi b}(0)$, completely determine the curve of length $L$, in three dimensional space.
@@
In three-dimensional Euclidean space, given any arbitrary but continuous and differentiable curve ${\bi r}(s)$, the derivatives $d{\bi r}(s)/ds$, $d^{2}{\bi r}(s)/ds^{2}$ and $d^{3}{\bi r}(s)/ds^{3}$, will be in general, three linearly independent vectors. If we take the next order derivative $d^{4}{\bi r}(s)/ds^{4}$, it can be expressed as a linear combination of the other three derivatives where the coefficients are functions of the four derivatives. This means that in the Frenet-Serret formalism the most general differential equation satisfied by a continuous and differentiable geometric curve in three dimensional space is a system of differential equations of fourth order where the independent variable is the arc length $s$.
@@
Let us go one step forward. Instead of doing Geometry let us make Physics by introducing clocks. Let us consider that we describe the electron by the time evolution in three-dimensional space of a point ${\bi r}(s)$. In the above geometrical analysis it will mean that there exists a continuous and differentiable function $s(t)$ between the arc length $s$ and the elapsed time $t$, that will represent from now on the parameter of the curve ${\bi r}(t)\equiv{\bi r}(s(t))$. Let us consider that an inertial observer describes the free motion of this electron. Because in general, all derivatives are different from zero, the motion of the point is accelerated and therefore this point ${\bi r}(t)$ cannot represent the location of the center of mass (CM) of the electron. It can represent perhaps the location of another property of the electron.
@@
If we are able to obtain a Lagrangian description of this system of only three degrees of freedom, the free Lagrangian $L_{0}$ has to depend up to the second time derivative of the point $\bi r$}, to obtain from the Euler-Lagrange equations a system of differential equations of fourth order. If this electron interacts there will be an interaction Lagrangian $L_{I}$, function of the interacting parameter, the charge, and the point $\bi r$} and some of its time derivatives. This is telling us that the point $\bi r$} represents the point where the external force is defined, i.e., the center of charge (CC) of the electron. This point has to be a different point than the possible CM, because in the free motion the CM must describe a straight line with no acceleration.
@@
Let us consider that in an inertial reference frame the motion is free. This means that for this inertial observer the displacement of the point at any time $t$, $ds(t)$ must be independent of the time. Otherwise if at two different times $t_{1}$ and $t_{2}$, $ds(t_{1})\neq ds(t_{2})$ this means that something different is happening at time $t_{2}$ than at time $t_{1}$, that produces a different displacement, contradictory with the idea that the motion is free. For this inertial observer the free motion of the electron is such that the CC is moving at a velocity of constant absolute value. If this description has to be independent of the inertial observer we arrive to the conclusion that we can never find an inertial observer at rest with respect to this point. If at some instant $t$, an inertial observer is at rest with respect to the point ${\bi r}(t)$, the velocity of this point is 0, but because this point is accelerated, at time $t+dt$ the velocity will be different from zero, contradictory with the condition that the velocity is of constant absolute value in any inertial frame. The velocity of the CC of an electron described by a single point must necessarily be an unreachable velocity of constant absolute value for all inertial observers.
@@
For the free motion also $\kappa(t)$ and $\tau(t)$, or its ratio $\kappa(t)/\tau(t)$ must be independent of the time, so that the trajectory of the CC is a helix. The center of the helix is a good candidate for the location of the CM that in the free motion is moving along a straight line at a constant velocity that can never reach the velocity of the CC. But the location of this point and the angular velocity of the comoving Frenet-Serret frame are completely determined by the point $\bi r$} and its subsequent time derivatives. The comoving frame rotates and the CC moves around the CM and because of this rotation and the orbital relative motion, this object has angular momentum. The CC has a trembling motion (zitterbewegung) around the CM. The particle described in this way has an internal frequency associated to this relative motion.
@@
2 General description and conclusions
The previous introduction is basically the Preamble of the Lecture Notes [1]. However, from a historical point of view, we arrived at this description of the electron by a different path. In 1985 [2] we analyzed the Lagrangian description of a mechanical system such that the boundary variables manifold that define the initial and final states of its variational evolution were considered to be a homogeneous space of the Galilei group. With this statement many models of non-relativistic spinning particles were described. The relativistic version of the above models were obtained in 1989 [3] by assuming that the boundary variables manifold of the variational formalism were homogeneous spaces of the Poincaré group. In that article we also assumed that the velocity of the point $\bi r$} could be the speed of light. Different models of relativistic spinning particles were found and also a spinning model of a classical photon described as a point and a comoving cartesian frame that rotates along the direction of the velocity. The spin of the photon has the same direction than the angular velocity, but it is an invariant mechanical property. The angular velocity and thus, the frequency of the photon, transform according to the Doppler effect transformation while the absolute value of the spin remains invariant.
@@ 2 General description and conclusions
The challenge is to find a definition based on group theory within the classical Lagrangian formalism. If we select any initial state of the variational description and transform this state by any Poincaré transformation we can obtain all possible physical states equivalent to the first one. If the mechanical system has no more classical states than the ones obtained by the above method, then, the set of states, i.e., the boundary variables manifold, is a homogenous space of the Poincaré group. If the mechanical system is not an elementary particle and we consider an excited state, we can never reach that state from the ground state by a simple change of the reference frame. The set of states of a non-elementary mechanical system is not a homogenenous space of the Poincaré group.
@@
The idea that leads to the definition of a classical elementary particle was revisited in 2008 in the form of a new fundamental principle, the Atomic principle [7]. The definition of elementary particle represents the physical fact that the elementary particle has no excited states and if it is not anihilated by some interaction its internal structure cannot be modified. A bound system like an atom can have excited states but an elementary particle cannot. Any mechanical system that contains some internal parts can be deformed but an elementary particle cannot. An elementary particle cannot be deformed and therefore all its possible states are just kinematical modifications of any one of them. In this way the whole classical theory of elementary spinning particles is based on three fundamental principles:
@@
2. The Variational Principle, postulates the existence of a Lagrangian that describes the elementary particle such that the dynamical equations are the Euler-Lagrange equations, and with the above principle they must be Poincaré invariant.
@@
3. The atomic Principle, i.e., the above definition of elementary particle. If the initial state in the Lagrangian description in an inertial reference frame is $x\equiv(x_{1},\ldots,x_{n})$, described by an unknown set of $n$ classical variables and this state changes at a later time to another $y\equiv(y_{1},\ldots,y_{n})$, and if the internal structure of the elementary particle has not been modified, then it is possible to find at this later time, another inertial observer, related to the previous one by means a Poincaré transformation $g$, such that the variables $y$ for this observer take exactly the same values as the variables $x$ for the previous observer. The the state $x$ is transformed into the $y$, such that the state y=gxy=g*x or $x=g^{-1}*y$. And this has to be valid for any pair of states. We recover that the boundary variables manifold of the Lagrangian description is a homogenous space of the Poincaré group.
@@
The quantum mechanical theory of elementary particles is also based on three fundamental principles. We have to substitute the Variational Principle by Feynman’s Quantization Principle such that the Hilbert space that describes the quantum mechanical states is such that any wave function is a complex, squared integrable function defined on the boundary variables manifold of the corresponding classical Lagrangian description [5]-[6]. With this quantization method, under an arbitrary Poincaré transformation, the wave function $\psi(x)$ transforms with a unitary irreducible representation of the Poincaré group if the point $x$ belongs to a homogeneous space of the group. For other groups, like the Galilei group, what is obtained is that the wave function transforms with a projective unitary irreducible representation of the group, i.e., a unitary representation up to an arbitrary phase. This is a technical feature that depends on the structure of the group. If the group has nontrivial central extensions, like the Galilei group, then the representation is not a true representation but a projective representation. The Poincaré group has only trivial central extensions and thus all representations are true representations. In the Lagrangian formalism, Lagrangians are Poincaré invariant but not Galilei invariant. They transform with a time derivative of a gauge function [14].
@@
The formalism can be extended to any kinematical group, Galilei, Poincaré, Weyl, Conformal or De Sitter groups, or any group we consider that describes the space-time symmetries of our mechanical system. If the rotation group is a subgroup of this general kinematical group we always obtain that the elementary particles described in this formalism have spin. If for instance, we select as the boundary variables manifold the space-time, i.e. the states are defined by the variables time $t$ and the position of a point $\bi r$}, $x\equiv(t,{\bi r})$, this manifold is in fact a homogeneous space of the Galilei or Poincaré group, and therefore it can be used to describe an elementary particle: The spinless point particle. But in Nature there are no spinless elementary particles. We need to include some more variables that transform under rotations, but we are restricted to a set that belongs to a homogeneous space of the kinematical group.

### v2

@@ Abstract
The answer to the title-question is affirmative. The analysis of the geometry of continuous and differentiable curves in three-dimensional Euclidean space suggests that the point represents the location of the center of charge of the electron, satisfies a system of ordinary differential equations of fourth order, and moves at the speed of light. The center of mass of the electron is a different point and will be determined by the evolution of the center of charge. It is the relative motion of the center of charge around the center of mass that gives rise to the spin and magnetic properties. The invariance of the mass and the absolute value of the spin for the center of mass observer imply that in the interaction of the electron with an external electromagnetic field the particle has to radiate.
@@ 1.1 Geometry
The theory of continuous and differentiable curves in three-dimensional space was developed in 1847 and is known as the Frenet-Serret formalism. The points of the curve are parameterized by the arc length $s$ that represents the Euclidean distance along the curve of this point ${\bi r}(s)$ from some initial position ${\bi r}(0)$. The function ${\bi r}(s)$ and its subsequent derivatives with respect to the parameter $s$ are continuous and differentiable functions of this parameter.
@@
Along the curve, and in terms of the derivatives of the point ${\bi r}(s)$ up to the third order, it is defined a system of three orthogonal dimensionless unit vectors attached to the point. They are called the tangent $\bi t$}, the normal $\bi n$} and the binormal $\bi b$}. The intrinsic characteristics of any curve are the two continuous and differentiable scalar functions $\kappa(s)$ and $\tau(s)$, known as the curvature and torsion, respectively, defined at every point ${\bi r}(s)$. The curvature is expressed in terms of the first two derivatives of the function ${\bi r}(s)$ and the torsion involves the derivatives up to the third order. The knowledge of the two continuous functions $\kappa(s)$ and $\tau(s)$, for $s\in[0,L]$, together the boundary conditions ${\bi r}(0)$ and ${\bi t}(0)$, ${\bi n}(0)$ and ${\bi b}(0)$, completely determine the curve of length $L$, in three dimensional space.
In three-dimensional Euclidean space, given any arbitrary but continuous and differentiable curve ${\bi r}(s)$, the derivatives $d{\bi r}(s)/ds$, $d^{2}{\bi r}(s)/ds^{2}$ and $d^{3}{\bi r}(s)/ds^{3}$, will be in general, three linearly independent vectors. The three Frenet-Serret unit vectors are written in terms of the derivatives up to the third order of point $\bi r$}, as:
${\bi t}=\frac{d{\bi r}}{ds},\quad{\bi n}=\frac{1}{\kappa}\frac{d^{2}{\bi r}}{ds^{2}},\quad{\bi b}=\frac{\kappa}{\tau}\frac{d{\bi r}}{ds}-\frac{\dot{\kappa}}{\kappa^{2}\tau}\frac{d^{2}{\bi r}}{ds^{2}}+\frac{1}{\kappa\tau}\frac{d^{3}{\bi r}}{ds^{3}},$ (1)
or conversely
$\frac{d{\bi r}}{ds}={\bi t},\quad\frac{d^{2}{\bi r}}{ds^{2}}=\kappa{\bi n},\quad\frac{d^{3}{\bi r}}{ds^{3}}=-\kappa^{2}{\bi t}+\dot{\kappa}{\bi n}+\kappa\tau{\bi b},$ (2)
where
$\kappa=\bigg|\frac{d^{2}{\bi r}}{ds^{2}}\bigg|,\qquad\tau=\frac{1}{\kappa^{2}}\left(\frac{d{\bi r}}{ds}\times\frac{d^{2}{\bi r}}{ds^{2}}\right)\cdot\frac{d^{3}{\bi r}}{ds^{3}},$ (3)
and have dimensions of length-1.
The derivatives of the unit vectors with respect to $s$, are expressed as:
$\dot{\bi t}=\boldsymbol{\omega}\times{\bi t}=\kappa{\bi n},\quad\dot{\bi n}=\boldsymbol{\omega}\times{\bi n}=-\kappa{\bi t}+\tau{\bi b},\quad\dot{\bi b}=\boldsymbol{\omega}\times{\bi b}=-\tau{\bi n},$ (4)
in terms of the Darboux vector $\boldsymbol{\omega}=\tau{\bi t}+\kappa{\bi b}$, with dimensions of length-1, where the dot represents the derivative with respect to $s$.
@@
If we take in (2) the next order derivative $d^{4}{\bi r}/ds^{4}$, taking into account (4) and (3) it can be expressed as a linear combination of the three unit vectors where the coefficients are functions of the four derivatives. These three unit vectors are expressed in (1) in terms of the derivatives of the point $\bi r$} up to the third order, and therefore we obtain a relationship between the derivatives of $\bi r$} up to order four. This means that in the Frenet-Serret formalism the most general differential equation satisfied by a continuous and differentiable geometric curve in three dimensional space is a system of ordinary differential equations of fourth order where the independent variable is the arc length $s$.
1.2 Physics
@@
Let us go one step forward. Instead of doing Geometry let us make Physics by introducing clocks. Let us consider that we describe the electron by the time evolution in three-dimensional space of a point $\bi r$}. In the above geometrical analysis it will mean that there exists a continuous and differentiable function $s(t)$ between the arc length $s$ and the elapsed time $t$, that will represent from now on the parameter of the curve ${\bi r}(t)\equiv{\bi r}(s(t))$. Let us consider that an inertial observer describes the free motion of this electron. Because in general, all derivatives are different from zero, the motion of the point is accelerated and therefore this point ${\bi r}(t)$ cannot represent the location of the center of mass (CM) of the electron. It can represent perhaps the location of another property of the electron.
@@
If we are able to obtain a Lagrangian description of this system of only three degrees of freedom, the free Lagrangian $L_{0}$ has to depend up to the second time derivative of the point $\bi r$}, to obtain from the Euler-Lagrange equations a system of differential equations of fourth order. If this electron interacts there will be an interaction Lagrangian $L_{I}$, function of the interacting parameter, the charge, the point $\bi r$} and some of its time derivatives and of external fields defined at point $\bi r$}. This is telling us that the point $\bi r$} represents the point where the external fields are evaluated. The external force is evaluated at this point, i.e., it represents the center of charge (CC) of the electron. This point has to be a different point than the possible CM, because in the free motion the CM must describe a straight line with no acceleration.
@@
Let us consider that in an inertial reference frame the motion of point $\bi r$}, although accelerated, represents a free motion. This means that for this inertial observer the displacement of the point at any time $t$, $ds(t)$, must be independent of the time. Otherwise if at two different times $t_{1}$ and $t_{2}$, $ds(t_{1})\neq ds(t_{2})$ this means that something different is happening at time $t_{2}$ than at time $t_{1}$, that produces a different displacement, contradictory with the idea that the motion is free. For this inertial observer the free motion of the center of charge of the electron is moving at a velocity of constant absolute value. If this description has to be independent of the inertial observer we arrive to the conclusion that we can never find an inertial observer at rest with respect to this point.
If at some instant $t$, an instantaneous inertial observer is at rest with respect to the point ${\bi r}(t)$, the velocity of this point is 0, but because this point is accelerated, at time $t+dt$ the velocity will be different from zero, contradictory with the condition that the velocity is of constant absolute value in any inertial frame.
The general motion of the point $\bi r$} has curvature and torsion. If it is moving at a velocity of constant absolute value for some inertial observer, the velocity is no longer of constant absolute value for another inertial observer moving at a constant velocity $\bi v$} with respect to the first one. If we add a constant vector $\bi v$} to a changing velocity vector does not produce a vector of a constant absolute value. There is nevertheless a unique solution if the velocities do not add according to the nonrelativistic linear addition. The velocity of the CC of an electron described by a single point must necessarily be an unreachable velocity of constant absolute value for all inertial observers. The velocity of the point $\bi r$} must transform according to the relativistic addition formula and it must be the speed of light. This is consistent with Dirac’s analysis of the electron such that the point $\bi r$} of Dirac’s spinor $\psi(t,{\bi r})$, where the external fields are evaluated, is in fact moving at the speed of light for all inertial observers. Dirac’s spinor represents the probability amplitude of the electron around the CC.
@@
For the free motion also $\kappa(t)$ and $\tau(t)$, or its ratio $\kappa(t)/\tau(t)$ must be independent of the time, so that the trajectory of the CC is a helix. The center of the helix is a good candidate for the location of the CM that in the free motion is moving along a straight line at a constant velocity that can never reach the velocity of the CC. But the location of this point and the angular velocity of the comoving Frenet-Serret frame are completely determined by the point $\bi r$} and its subsequent time derivatives. The comoving frame rotates and the CC moves around the CM and because of this rotation and the orbital relative motion, this object has angular momentum. The CC has a trembling motion (zitterbewegung) around the CM. It is not necessary to postulate, as in De Broglie formalism, that elementary matter has an internal frequency. The particle described in this way has an internal frequency associated to this relative motion.
@@
1.3 Invariance arguments
Let us assume that an inertial observer describes the evolution of the electron by the trajectory of a point ${\bi r}(t)$ in three-dimensional space. If we make an arbitrary Poincaré transformation to any other inertial observer what we get is the family of all trajectories ${\bi r}(t,g)$ of the point for all inertial observers, where $t$ is the time coordinate in each frame. This family is parameterized by the 10 parameters that characterize the arbitrary Poincaré transformation $g$. All these trajectories satisfy a Poincaré invariant differential equation obtained from the above family ${\bi r}(t,g)$ and its subsequent time derivatives, by eliminating the 10 parameters in terms of the different derivatives. To eliminate 10 parameters we need to reach up to the fourth-order derivative to have sufficient equations, and therefore we find again that the trajectory of the point for any inertial observer satisfies a system of Poincaré invariant ordinary fourth order differential equations. This dynamical equation is Poincaré invariant by construction because is independent of all group parameters.
The point is not the center of mass, and in the free motion, if it represents the center of charge it is accelerated but cannot radiate because energy is conserved. A particle with CC and CM as different points can never radiate if the CM is not accelerated. A free electron never radiates. The Lorentz-Dirac theory of radiation considers that both points are the same, but this is an unphysical constraint. The center of mass represents the location of the center of inertia of the particle while the CC represents the location of the center of interaction. Inertia and interaction are different physical properties of matter. Theoretical physics is oblidged to analyze physical systems where these two points could be different.
2 General description
The previous introduction is contained basically in the Preamble of the Lecture Notes [1]. However, from a historical point of view, we arrived at this description of the electron by a different path. In 1985 [2] we analyzed the Lagrangian description of a mechanical system such that the boundary variables manifold that define the initial and final states of its variational evolution were considered to be a homogeneous space of the Galilei group. With this statement many models of non-relativistic spinning particles were described. The relativistic version of the above models were obtained in 1989 [3] by assuming that the boundary variables manifold of the variational formalism were homogeneous spaces of the Poincaré group. In that article we also assumed that the velocity of the point $\bi r$} could be the speed of light, as suggested by Dirac’s analysis. Different models of relativistic spinning particles were found and also a spinning model of a classical photon described as a point and a comoving cartesian frame that rotates along the direction of the velocity. The spin of the photon has the same direction than the angular velocity, but it is an invariant mechanical property. The angular velocity and thus, the frequency of the photon, transform according to the Doppler effect transformation while the absolute value of the spin remains invariant.
@@ 2 General description
The challenge is to find a definition based on group theory within the classical Lagrangian formalism. If we select any initial state of the variational description and transform this state by any Poincaré transformation we can obtain all possible physical states equivalent to the first one. If the mechanical system has no more classical states than the ones obtained by the above method, then, the set of states, i.e., the boundary variables manifold, is a homogeneous space of the Poincaré group. If the mechanical system is not an elementary particle and we consider an excited state, we can never reach that state from the ground state by a simple change of the reference frame. The set of states of a non-elementary mechanical system is not a homogeneous space of the Poincaré group.
@@
The idea that leads to the definition of a classical elementary particle was revisited in 2008 in the form of a new fundamental principle, the Atomic principle [7]. The definition of elementary particle represents the physical fact that the elementary particle has no excited states and if it is not annihilated by some interaction its internal structure cannot be modified. A bound system like an atom can have excited states but an elementary particle cannot. Any mechanical system that contains some internal parts can be deformed but an elementary particle cannot. An elementary particle cannot be deformed and therefore all its possible states are just kinematical modifications of any one of them. In this way the whole classical theory of elementary spinning particles is based on three fundamental principles:
@@
2. The Variational Principle, postulates the existence of a Lagrangian that describes the elementary particle such that the dynamical equations are the Euler-Lagrange equations, and with the above principle they must be Poincaré invariant. But we do not know yet what are the independent degrees of freedom. This is determined by the next fundamental principle.
@@
3. The atomic Principle, i.e., the above definition of elementary particle. If the initial state in the Lagrangian description in an inertial reference frame is $x\equiv(x_{1},\ldots,x_{n})$, described by an unknown set of $n$ classical variables and this state changes at a later time to another $y\equiv(y_{1},\ldots,y_{n})$, and if the internal structure of the elementary particle has not been modified, then it is possible to find at this later time, another inertial observer, related to the previous one by means a Poincaré transformation $g$, such that the variables $y$ for this observer take exactly the same values as the variables $x$ for the previous observer. Then the state $x$ is transformed into the $y$, such that the state y=gxy=g*x or $x=g^{-1}*y$. And this has to be valid for any pair of states. We recover that the boundary variables manifold of the Lagrangian description is a homogenous space of the Poincaré group.
The quantum mechanical theory of elementary particles is also based on three fundamental principles. In the classical description, the Lagrangian formalism determines the path of the system between the initial point $x_{1}$ to the final point $x_{2}$. Because we are not going to measure the path followed by the system we assume Feynman’s hypothesis that all paths between these two points are allowed with the same probability. We have to substitute the Variational Principle by Feynman’s Quantization Principle such that the Hilbert space that describes the quantum mechanical states is such that any wave function is a complex, squared integrable function defined on the boundary variables manifold of the corresponding classical Lagrangian description [5]-[6].
@@
With this quantization method, the wave function $\psi(x)$ represents the probability amplitude of finding the electron around point $x$. Under an arbitrary Poincaré transformation $\psi(x)$ transforms with a unitary irreducible representation of the Poincaré group if the point $x$ belongs to a homogeneous space of the group. For other groups, like the Galilei group, what is obtained is that the wave function transforms with a projective unitary irreducible representation of the group, i.e., a unitary representation up to an arbitrary phase. This is a technical feature that depends on the structure of the group. If the group has nontrivial central extensions, like the Galilei group, then the representation is not a true representation but a projective representation. The Poincaré group has only trivial central extensions and thus all representations are true representations. In the Lagrangian formalism, Lagrangians are Poincaré invariant but not Galilei invariant. They transform with a time derivative of a gauge function [8].
@@
The formalism can be extended to any kinematical group, Galilei, Poincaré, Weyl, Conformal or De Sitter groups, or any group we consider that describes the space-time symmetries of our mechanical system. If the rotation group is a subgroup of this general kinematical group we always obtain that the elementary particles described in this formalism have spin. If for instance, we select as the boundary variables manifold the space-time, i.e. the states are defined by the variables time $t$ and the position of a point $\bi r$}, $x\equiv(t,{\bi r})$, this manifold is in fact a homogeneous space of the Galilei or Poincaré group, and therefore it can be used to describe an elementary particle: The spinless point particle. The angular momentum of this particle with respect to the point $\bi r$} is zero. But in Nature there are no spinless elementary particles. We need to include some more variables that transform under rotations, but we are restricted to a set, larger than space-time, that belongs to a homogeneous space of the kinematical group. In that case, the angular momentum of the particle with respect to the point $\bi r$} is different from zero and the particle has spin.
@@
In the quantum version, Dirac’s spinor is a squared integrable complex function of the variables $\psi(t,{\bi r},{\bi u},\boldsymbol{\alpha})$. The constraint $u=c$ implies that it is a function of the non-compact variables $(t,{\bi r})$ and of the following five compact variables: Two are the zenithal angle $\beta$ and azimuthal angle $\lambda$} of the velocity $\bi u$} of constant value $u=c$ and the other three angles that describe the orientation of the attached local cartesian frame. If we represent this local frame as a rotation matrix it is characterized by the rotated angle $\alpha$ and the zenithal angle $\theta$ and azimuthal angle $\phi$} of the unit vector along of the rotation axis. The spin operators are linear differential operators with respect to these 5 compact variables and commute with the Hamiltonian $H=i\hbar\partial/\partial t$ and the linear momentum operator ${\bi p}=-i\hbar\nabla$.